OpenStudy (anonymous):

What is the slope of the line y= -1/4x - 3/4 A. -4 B. - 1/4 C. -3 D. -3/4

OpenStudy (the_fizicx99):

In y = mx + b The slope is M The y intercept is B

OpenStudy (the_fizicx99):

So in your formula, y= -1/4x - 3/4 What's the slope

OpenStudy (shaik0124):

b is slope

OpenStudy (anonymous):

no m is slope

OpenStudy (shaik0124):

-1/4 is answer

OpenStudy (the_fizicx99):

-.- Really??

OpenStudy (shaik0124):

i mean option B is answer

OpenStudy (the_fizicx99):

You just gave out the answer and interrupted meh? Smh ..

OpenStudy (shaik0124):

compare y=mx+c with y=-1/4x-3/4 here m=-1/4

OpenStudy (the_fizicx99):

You're not explaining why though.